OpenStudy (anonymous):

Find the integral of x/(sqrt(1-x^2)) dx.

OpenStudy (ybarrap):

$$ \int \cfrac{x}{\sqrt{1-x^2}} dx $$ Let \(u=1-x^2\) then \(du=-2xdx\): $$ \int \cfrac{-du}{2\sqrt{u}} dx=-\cfrac{1}{2}\int \cfrac{du}{\sqrt{u}} dx $$ Can you take it from here?

OpenStudy (anonymous):

Please wait.

OpenStudy (anonymous):

I got it. Thanks.

OpenStudy (ybarrap):

you're welcome